Please help me answer this question


Find the height of a triangle of area 70 square inches if its base is 20 in. long. State your answer in inches.

Answer by mananth(16946)

Find the height of a triangle of area 70 square inches if its base is 20 in. long. State your answer in inches.
...
Area of traingle = 1/2 * base * height
70=1/2*20*h
h= 2*70/20
h = 7 inches
..
m.ananth@hotmail.ca